A rectangular painting with a width of x centimeters has an area of x^2 + 50x square centimeters. find a bionominal that represents the length.
Answer by checkley77(12844)

AREA=LW
X^2 + 50X
X(X+50)
AREA=X(X+50)
AREA=W*L
W=X
L=(X+50) ANS.